You own 1,800 shares of stock in Avondale Corporation. You will receive a $1.80 per share dividend in one year. In two years, the company will pay a liquidating dividend of $80 per share. The required return on the company's stock is 25 percent. Suppose you want only $600 total in dividends the first year.

  

What will your homemade dividend be in two years? (Do not round intermediate calculations and round your answer to the nearest whole dollar amount, e.g., 32..)
Answer #1

Price of the stock in one year = $80 / 1.25 = $64.00

If you only want $600 in Year 1, you will buy:

($3,240 – $600) / $64.00 = 41.25 shares

at Year 1. Your dividend payment in Year 2 will be:

Year 2 dividend = (1,800 + 41.25)($80) = $147,300